18. You are comparing stock A to stock B. stocks should you prefer and why? ven the following information, which one of these two REXEL UNIVERSITY- State of Rate of Return if Probability of Economy State of Economy StoskA Stok A. Stock A; because it has B, C. Stock Stock A; because it has higher expected an expected return of 7% and appears to be more risky expected return and appears to be less risky than stock B because it has a slightly lower expected return but appears to stock B. D. Stock B; because i be significantly less risky than 4 t has a higher expected returm and appears to be just slightly more risky than as a higher expected return and appears to be less risky than stock A stock A. 19. Zelo, Inc. stock has a beta of 1.23. The i 10%, what is the amount of the risk premium on Zelo stock .23. The risk-free rate of return is 4.5% and the market rate of return is A. 4.47% B. 5.50% C. 5.54% D. 6.77% E. 12.30% 20. Which of the following statements is true? A. A well-diversified portfolio has negligible systematic risk. B. A well-diversified portfolio has negligible unsystematic r C. An individual security has negligible systematic risk. Anl security has negligible unsystematie E. Both A and D. stock pays a quarterly dividend of $.10 a share. Today, you sold all of your shares for S45.13 p What is the total amount of your capital gains on this investment? A. $1.24 21. Six months ago, you purchased 100 shares of stock in ABC Co. at a price of $43.89 a sha B. $1.64 C. $40.00 D. $124.00 E. $164.00 22. One year ago, you bought a stock at a price of $32 a share. Today, you sold the sto total return of 14.62%. Your capital gain was $3.48 a share. What was your dividend yi A. 2.25% B. 3.75% C. 3.35% D. 2.85% E. 4.35%

tHANK YOU
Answer #1
18) Option B
19) Amount of risk premium = 1.23*(10-4.5) = 6.77
Answer: Option D
20) Option A
21) (45.13-43.89)*100 = $124
Option: D
22) Capital gain yield = 3.48/32= 10.88%
Dividend yield = 14.62-10.88 = 3.74
Option: B (Closest answer)